OpenStudy (anonymous):

How do I find the common ratio with this data? Ball 1 height - 8ft, 5 1/2ft, 4ft, 2ft, 3/4ft Ball 2 height - 8ft, 5 3/4ft, 4 1/3ft, 2 1/2ft, and 1 ft Ball 3 height - 8ft, 5 1/3 ft, 4 1/4ft, 2 1/3ft, 1 ft

OpenStudy (anonymous):

@antoni7

OpenStudy (anonymous):

PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

generally the common ratio is found by : the first term divided by the second term

OpenStudy (anonymous):

you got no idea how to do this ?

OpenStudy (anonymous):

So I divide 5 1/2 by 4 ft?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

and i divide 2ft and 3/4ft?

OpenStudy (anonymous):

sorry if this is an obvious question...

OpenStudy (anonymous):

ball 1 height : common ratio = 8/ 5 1/2 ball 2 height : common ratio = 8 / 5 3/4 ball 3 height : common ration = 8 / 5 1/3

OpenStudy (anonymous):

do i divide those?

OpenStudy (anonymous):

you ll obtain the common ratio of each balls height...

OpenStudy (anonymous):

so is that a yes?

OpenStudy (anonymous):

yeah... what im confused about is the question ... does it ask you for each ball height common ratio ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i found my answers by dividing

OpenStudy (anonymous):

common ratio is usually denoted by " r " So r for ball height 1 = 16/5 height 2 = 32/15 height 3 = 24 /5

OpenStudy (anonymous):

so do i divide those?

OpenStudy (anonymous):

i divided the ones you showed me previously

OpenStudy (anonymous):

yup i just wrote the answers..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

just inform me if you got these answers right or wrong

OpenStudy (anonymous):

I definitely will.

OpenStudy (anonymous):

ok good luck :) Enjoy studying

OpenStudy (anonymous):

Do you think the common ratio would affect the size of the ball?

OpenStudy (anonymous):

I don't see how it would...?

OpenStudy (anonymous):

For Geometric sequences suppose you got a sequence : 2, 4, 16 , 64 the first term would be denoted as " a" and where a = 2 the common ratio , r = 4 divide 2 = 2 the general form , Tn = ar^ (n-1) Tn = 2 x 2 ^ ( n-1) suppose you want to know the 3rd term of the sequence T3 = 4 ^ (3-1) = 4 ^ 2 = 16

OpenStudy (anonymous):

so yes....?

OpenStudy (anonymous):

im confused by your example

OpenStudy (anonymous):

the common ratio is something which is constant..

OpenStudy (anonymous):

for a specific sequence the common ratio stay the same..

OpenStudy (anonymous):

oh alright:) wow thanks

OpenStudy (anonymous):

can you help me go through this other step please

OpenStudy (anonymous):

sure but i'm not 100% right about the answers i'm giving you... you need to check it out later with your teacher ....

OpenStudy (anonymous):

If ball 1 were dropped from 2 feet higher, would the common ratio be different? Explain your answer. I think it would because you'd be adding 2 feet more to the equation

OpenStudy (anonymous):

yeah i guess so

OpenStudy (anonymous):

10 ft divide by 8 common ratio , r = 10 / 8 = 1.25

OpenStudy (anonymous):

wheareas 8 ft divide by 5.5 ft , r = 8 / 5.5 = 1.45

OpenStudy (anonymous):

What is the height of each ball on the fifth bounce (i.e., Height 6)? Use the geometric sequence formula, an = a1rn – 1 and show your work. an is the last term a1 = 1st term r = ratio n = # of terms in a finite sequence

OpenStudy (anonymous):

What numbers would I use to plug in?

OpenStudy (anonymous):

there were 3 balls

OpenStudy (anonymous):

T6 = first term x (common ratio of ball 1height )^ ( 5)

OpenStudy (anonymous):

T6 = first term x (common ratio of ball 2height )^ ( 5) T6 = first term x (common ratio of ball 3height )^ ( 5)

OpenStudy (anonymous):

thank you.... oh my god thank you thank you thank you

OpenStudy (anonymous):

where do i find the 1st terms though?

OpenStudy (anonymous):

one quick question

OpenStudy (anonymous):

where do i find the ball height? are they talking about the very first one in the data i provided?

OpenStudy (anonymous):

first term = 8 ft

OpenStudy (anonymous):

8ft x 3.2 x ___________?

OpenStudy (anonymous):

a1 x ratio x height of 1^5

OpenStudy (anonymous):

whats the height of 1, 2, and 3?

OpenStudy (anonymous):

its the first term of each ball that is ball 1 = 8ft ball 2= 8ft ball 3 = 8ft

OpenStudy (anonymous):

wait huh?

OpenStudy (anonymous):

i just use 1 for the first height, 2 for the second height, and 3 for the third?

OpenStudy (anonymous):

wait wait "confused" what is the question ?

OpenStudy (anonymous):

what I use is: a1 x ratio x "ball height" 1^5 but what do i plug in for "ball height"?

OpenStudy (anonymous):

the ball height is 8ft right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that's the first term (a1)

OpenStudy (anonymous):

T6 = first term x (common ratio of ball 1height )^ ( 5) = (8 x 3.2 )^5 T6 = first term x (common ratio of ball 2height )^ ( 5) = (8 x 2.13333 )^5 T6 = first term x (common ratio of ball 3height )^ ( 5) = (8 x 4.8 ) ^ 5

OpenStudy (anonymous):

its the answer for the question : What is the height of each ball on the fifth bounce (i.e., Height 6)? Use the geometric sequence formula, an = a1rn – 1 and show your work. an is the last term a1 = 1st term r = ratio n = # of terms in a finite sequence

OpenStudy (anonymous):

ohhhh!!!!!!!!!!!!!! so what threw me off was the height

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

done with this awesome question on geometric series ?

OpenStudy (anonymous):

yes thank god

OpenStudy (anonymous):

but i literally have one more question that deals with all of this

OpenStudy (anonymous):

I have to find the total distance of the height each ball was traveled using a specific formula

OpenStudy (anonymous):